OpenStudy (anonymous):

what is the value of x 14-x=21 x=

OpenStudy (michele_laino):

Please try to subtract 14 to both sides of your equation first

OpenStudy (anonymous):

0 and 7 @Michele_Laino

OpenStudy (michele_laino):

you should get this: -x=7

OpenStudy (anonymous):

yea i got that @Michele_Laino

OpenStudy (michele_laino):

ok! Now multiply both sides by -1

OpenStudy (anonymous):

14 and 21 @Michele_Laino

OpenStudy (michele_laino):

Sorry I wanted say this: multiply both sides of the equation -x=7, by -1

OpenStudy (anonymous):

do u know the answer ? @Michele_Laino important

OpenStudy (michele_laino):

hint: -1*(-x)=x, and -1*7=-7 so what is the solution?

OpenStudy (anonymous):

7? @Michele_Laino

OpenStudy (michele_laino):

no, if -1*7=-7, then?

OpenStudy (anonymous):

-7 @Michele_Laino

OpenStudy (michele_laino):

ok! That's right!

OpenStudy (anonymous):

thanks! @Michele_Laino

OpenStudy (michele_laino):

thanks! @mamas

OpenStudy (anonymous):

can u help me with the rest ? @Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

-5x=-35 what is the value of x ?

OpenStudy (michele_laino):

please, multiply both sides of that equation, by -1, first

OpenStudy (anonymous):

-6 and -36 ? @Michele_Laino ?

OpenStudy (michele_laino):

no, hint: -1*(-35)=35 and: -1*(-5x)=5x

OpenStudy (michele_laino):

so we have: 5x=35

OpenStudy (anonymous):

-5 ?

OpenStudy (michele_laino):

please divide both sides of the equation 5x=35 by 5

OpenStudy (lizz123):

5x=35 which number has an x next to it

OpenStudy (anonymous):

5 @lizz123

OpenStudy (lizz123):

k you divide 5 by 5x and 35

OpenStudy (lizz123):